OpenStudy (anonymous):

Check my work! :) Can \(2y^2 - 5y = -2\) become \(2y^2 - 5y + -2 = 0\)?

OpenStudy (dan815):

the change is incorrect, it should be +2

OpenStudy (anonymous):

Hello, Dan. :)

OpenStudy (dan815):

because you are going to add 2 to both sides to eliminate the -2 on the right

OpenStudy (anonymous):

Right. Sorry, I forgot. Thank you for the correction.

OpenStudy (dan815):

Okay, now get to factoring

OpenStudy (anonymous):

I am learning the Quadratic Formula and I had to get my equation to look like this: \[-b \pm \sqrt{b^2 - 4ac} \over 2a\]

OpenStudy (dan815):

\[2y^2-5y+2=0\\\] a=2, -5=b, 2=c

OpenStudy (dan815):

you see the coefficients, a is the coefficient on the y^2(other wise known as the 2nd degree term), b is the coefficient on the y term( of the 1st degree term) and c is the constant ( or the 0th degree term)

OpenStudy (anonymous):

Yes, I see. :)